# Probes of Anomalous Events at LHC with Self-Organizing Maps

### v2

@@ Abstract
We propose an Unsupervised Learning Algorithm, Self-Organizing Maps (SOM), built on a neural network architecture, for the probe of a rare top decay, mediated by Flavor Changing Neutral Current (FCNC), to charm and the Higgs boson, with the Higgs boson further decaying to a pair of b-quarks or a pair of gauge bosons ($W^{\pm}/Z$) in a boosted regime. Ideally, the particles originating from the decay of the boosted top lead to the reconstruction of a large-R jet, comprising three-prong substructures, with b- and c-tagged subjets. We have illustrated that SOM can be effectively implemented as an anomaly-finder for model-agnostic probes of such decay at the LHC. Relatively simple in formalism, this algorithm achieves commendable performance over other existing formalisms, tailored to probe this rare decay. We discuss the robustness of this algorithm, which can be successfully implemented for other BSM probes, both for model-agnostic and model-dependent searches.
@@ 1 Introduction
As an illustration, we probed a rare top decay, driven by a flavor changing neutral current (FCNC), to a c-quark and Higgs boson $\left(t\to cH\right)$ with suppressed branching fraction $\lesssim 0.1\%$ [43, 44] in a boosted regime ($p^{top}_{T}>$ 350 GeV), over the background dominated by the most abundant decay mode of the top $\left(t\to bW\right)$, and spurious QCD events $\left(pp\to jj\right)$ mimicking the signal topology. The decay products of the boosted top are highly collimated, enveloped by a large-R jet comprising a three-prong substructure at LHC. Kinematic features specific to the signal $\left(t\to cH\right)$, irreducible background $\left(t\to bW\right)$, and the QCD di-jet background, are mapped on the planar grid which, through iterative adjustment of weights attributed on the nodes, leads to the identification of individual decay modes as clusters appearing on the grids. This probe of rare top decay mediated by FCNC using SOM achieves a signal significance of $\sim 2\sigma$ for $\mathcal{BR}(t\to cH)$ = 0.05% at 3 ab-1 over SM backgrounds, a fair leap as compared to our previous endeavor [45], where we presented a simple ML model for the same probe that outperformed existing top-tagging methodologies [40], tailored to probe the said decay. Finally, we provide an interpretation for each cluster boundary formed on the SOM grid, drawing insights into the relevant features underlying the event topologies of the three different processes under consideration.
@@ 2 Self Organizing Maps: The Algorithm
where, $v$ are the coordinates of the said node, $u$ are the coordinates of the BMU, $W_{v}(t)$ is the weight vector of the said node at time step $t$, $X(t)$ is the input feature presented at time step $t$, $\eta(t)$ is time-dependent learning rate, and $\phi(u,v,t)$ is the Neighborhood Distance Function for the BMU (at $u$) and the concerned node (at $v$) at the time step $t$. The learning rate $\eta(t)$ follows an exponential damping given as,
@@ 3 Event Generation and Reconstructed Objects
      N-subjettiness ($\tau^{\beta}_{N}$): N-subjettiness characterizes the substructure of boosted jets [61]. Considering N hypothetical subjet axes, N-subjettiness is determined by summing up the angular distances of jet constituents with respect to their respective closest subjet axes. Essentially, it indicates how well a jet aligns with N subjet-structure hypothesis. To establish the prominence of the existence of three subjet axes within the large-R jets of the signal (or the irreducible background) over two, and two subjet axes over one (for QCD di-jet background), we analyze the ratios of N-subjettiness measurements, specifically measuring $\tau_{N}^{\beta}/\tau_{N-1}^{\beta}$, labeled $\tau_{N(N-1)}^{\beta}$, for respectively $N=3,~{}2$, considering the angular weightage exponent $\beta=1$. These ratios enables the isolation of the QCD multijets from top jet samples.
@@
      Energy- and Mass-Fraction Information: The energy repatriation of the large-R jet is evaluated by the fraction of its total energy stored in its constituent subjets. For a given subjet, this is measured as, $\Delta E_{T}^{subjet}=\frac{E_{subjet}}{E_{topjet}}$. We further estimate the composition of large-R jet, with an observable measuring the fractional mass retained after removing the invariant mass of specific subjet combinations (indexed by $i,~{}j$): $\Delta X_{ij}=\left(1-\frac{M_{ij}}{M_{topjet}}\right)$. Accurate identification of light jets is crucial for this analysis, and is achieved through a $\chi^{2}$-analysis detailed in [45].
@@ 4.2 Robustness of the cluster boundaries
    where, ‘$a$’ denotes the mean distance between e and all other events within its own cluster, and ‘$b$’ is the mean distance between e and all events in its nearest-neighboring cluster. The average value of $S$ for all events gives the average Silhouette score. This score ranges between $[-1,1]$, where positive values indicate moderate to good clustering and negative values suggest poor clustering. A score close to zero implies that the events are globally populated in the vicinity of the cluster boundaries.
@@ 4.4 Clustering efficiency with Test dataset
    Test dataset2A/2B/2C: $pp\to tW$ events with top quark decaying through an FCNC decay, $t\to cH$ followed by $H\to b\bar{b}$/$H\to WW^{*}$/$H\to ZZ^{*}$ decay. The $W$-boson produced in association with the top quark decays leptonically.
@@
    Test dataset3A/3B: The most dominant channel, $pp\to t\bar{t}$, with one of the tops decaying leptonically111 The simulation methodology for the $t\bar{t}$ events, including the software suites, is identical to that described in Section 3. $(t\to bW,W\to\ell\nu_{\ell})$, and the other top undergoes $t\to bW$ decay or $t\to cH$ decay with $W$/$H$ decaying hadronically.
@@ 5 BSM Reach: Event Selection to Signal Significances
[…]
For those events which pass the pre-selection cuts, we identify three signal regions (SR) based on the flavor tagging information among the subjets of the large-R jet. The signal regions are primarily motivated by the understanding of the different clusters obtained after training the SOM and their characteristics in terms of the high-level objects. Additionally, for a comparative analysis, the SRs are also defined keeping in mind the resolved (non-boosted) analysis performed by the ATLAS Collaboration to find the top FCNC decay at the 13 TeV run of LHC.
@@
The first signal region (SR-I) is almost the same as the ATLAS study, with the only difference being that the requirements are made at the subjet level. We extend the selection strategy of SR-I to SR-II, accommodating the further requirement of c-tagged jets within the large-R jet. It is needless to say that SR-II is expected to be the most efficient selection strategy in probing the top FCNC decay we are interested in. Finally, we also study the sensitivity of a relatively relaxed criterion to the multiplicity of b-tagged jets within the large R jet by introducing the third signal region, SR-III. Note that, for both SR-I and SR-II, the presence of at least two b-tagged jets enables us to calculate the invariant mass of the two b-tagged jets, and compare the reconstructed mass with the Higgs boson mass. In Figure 7, we show the distribution of the invariant mass mentioned above for the SM Top samples, Test dataset2A, Test dataset3A and Test dataset3B that satisfies the criteria for SR-I and SR-II without the invariant mass cut. For SR-I and SR-II, therefore, we further demand that the invariant mass of the two leading b-tagged jets satisfy the following: $M_{bb}=[110,140]~{}{\rm GeV}$. The details of the signal regions are listed in Table 4.
@@
Table 4: Signal regions defined to probe the top FCNC decay. Here the observables $nSubjet$, $N_{b-jet}$ and $N_{c-jet}$ denote the total number of subjets, number of b-tagged jets and number of c-tagged jets within the large-R jet, respectively. The invariant mass of the two leading b-tagged jets is denoted by the quantity $M_{bb}$.
As noted previously, the FCNC interactions between the top quark and the Higgs boson via the $t-c-H$ coupling are essential in our quest for BSM physics. To evaluate the feasibility of detecting this exotic mode in the boosted regime using the SOM algorithm, we compute the signal significance, defined as $\mathcal{S}=\sqrt{2\left[(S+B)\ln\left(1+\frac{S}{B}\right)-S\right]}$ [65] where S and B represent the number of signal and background events that meet the jet selection criteria and satisfy the trigger/pre-selection requirement. Before we proceed to calculate the signal significances for the three signal regions, it is important to estimate the cluster efficiencies for each of the SRs. In Table 5, we show the leading order (in $\alpha_{s}$) production cross section for all the processes considered here with $\sqrt{s}=13~{}{\rm TeV}$. The last column in Table 5 estimates the expected number of events at the 13 TeV run of LHC with integrated luminosity $\mathcal{L}=3000~{}fb^{-1}$. It is important to note that even though QCD di-jet events seem to overshadow signal events as seen in $3^{rd}$ column of $2^{nd}$ row, the number of events that eventually satisfy the trigger and jet selection cuts, and contribute to the signal regions defined in Table 4 is negligible.
Process Cross-section ($fb$) $\mathcal{BR}$ $\times\mathcal{BR}\times\mathcal{L}$
SM top 435 $0.11\times 1.0\times 0.68$ 9.76×$\times 10^{4}$
QCD di-jet $3.91\times 10^{6}$ 1 1.17×$\times 10^{10}$
Test dataset2A 435 $0.11\times 1.0\times 0.6$ 8.61×$\times 10^{4}$
@@ C-1 C-2 C-3 C-4 C-5 C-6 C-7
[…]
@@ 6 Summary and Outlook
Probes of processes involving rare decays driven by Flavor-Changing Neutral Current (FCNC) offer a powerful avenue for exploring Physics beyond the Standard Model. The rare decay, $t\to cH$, considered in this article, offers sensitive probe of potential New Physics interactions, in particular, the $t-c-H$ coupling. In contrast to the supervised learning methodology presented in [45], we present a complementary, model-independent analysis based on unsupervised learning technique, targeting the boosted regime of this decay. With judicious choices of jet-substructure based observables along with efficient flavor tagging of subjets, similar to the approach followed in [45], we achieve an enhanced signal significance over our previous analysis [45], which outperformed the performances of advanced ML-driven approaches, viz., GNN-based LorentzNet and Transformer-based classifiers.
@@
In [45], we argued that the judicious choice of jet-substructure based observables provided an edge over the GNN-based models which train on the 4-vectors of the large-R jet constituents. With the jet substructures of the signal $\left(t\to cH,~{}H\to b\bar{b}\right)$, bearing close resemblance to those of the irreducible background $\left(t\to bW,~{}W\to c\bar{s}\right)$, this schema achieved relatively lower significance. The transformer-based approaches implement attention mechanism, which determines the relevance of input features for proper identification signals from the background, thereby allowing a dense deep-learning classifier selectively focus on relevant features suppressing the irrelevant ones, consequently enhancing the accuracy and efficiency of the signal recognition. We observed that a simple boosting algorithm with decision tree stumps trained over judiciously chosen jet-substructure-based features, nevertheless, exhibited higher efficiency. This indicates that our choice of input features made the feature space so fragmented that the dominance of a specific class (signal or different SM backgrounds) in a given subspace becomes prominent.
@@
The Self Organizing Maps, on the contrary, follows a different approach. The nodes, placed at the intersection of its grids, attempt to mimic the feature vector for each input data-point (event topologies) iteratively. The weights are adjusted so that the best matching unit (BMU) and the nodes in its neighborhood have their respective weights aligned with the matched feature vector. With training over each iteration (and updating weights assigned to each node), and subsequently over every batch (and subsequent updation of model parameters), the model captures the latent features specific to signal and different backgrounds. Finally, the grids end up forming clusters, with every node in each cluster possessing a weight vector resembling the underlying features that distinguish different event classes. As discussed in Section 5, each cluster predominantly captures events belonging to a specific Test dataset (see, Tables 6 to 8). The clusters formed on SOM grid are the planar projection of the feature space of the events, in a way, that allows events with similar underlying topology captured in a single cluster.
@@
Training SOM with SM top quark and QCD di-jet samples with assessment of topographic error in clustering, we observe that the emerged map fairly preserves the data topography within the grid. Further testing of the trained model with FCNC-driven top decay to a charm quark and a Higgs boson ($t\to cH$) demonstrates remarkable performance in identification of the anomalous regions for the training samples. We define three signal regions (SRs) based on the three-prong substructure of the candidate top jet, and their kinematics. For SR-II, our model achieves a significance of $\sim 2\sigma$ for $\mathcal{BR}(t\to cH)$ = 0.05%, establishing the model performance for segregating rare FCNC-mediated decays from overwhelming SM backgrounds. We emphasize that a possible explanation behind an enhanced performance of SOM over the existing approaches is embedded in the working principle of the algorithm, which learns to mimic the event features, leading to better capturing of patterns specific to the signal and different backgrounds in distinct clusters.
We have considered the implementation of SOM as an unsupervised anomaly detector in Section 4, achieving fairly substantial significance despite vanishingly small $t\to cH$ branching fraction over an overwhelming SM background. We commenced with training the SOM architecture with SM backgrounds and used the trained model to probe the rare top-decay as anomalous events over the identified SM patterns. SOM can also be tailored for supervised learning, upon training the grids with the signal events and the SM backgrounds, thereby estimating the clusters and the event density in each cluster, using SOM as a model-dependent probe of the rare top-decay.
@@
SOMs, since its inception, has outgrown as tool for clustering and dimensionality reduction. As evoked above, it has been adapted for supervised learning, enabling enhanced model performance, simultaneously preserving the topography of the training dataset. This versatility makes SOM a potential tool for BSM probes. Further improvements can be explored with alternative distance metrics, viz., the Mahalanobis distance, or the cosine similarity, in order to obtain more distinct clusters of signals and backgrounds on Kohonen plane. Adaptive frameworks, such as, Growing Self-Organizing Map (GSOM), which dynamically adapts its structure during training, by adding nodes as required based on data complexity, could further optimize the signal-to-background ratio. SOMs and its extensions are therefore, potential tools for future phenomenological explorations.

### v3

@@ Abstract
We propose an Unsupervised Learning Algorithm, Self-Organizing Maps (SOM), built on a neural network architecture, for the probe of a rare top decay, mediated by Flavor Changing Neutral Current (FCNC), to charm and the Higgs boson, with the Higgs boson further decaying to a pair of b-quarks or a pair of gauge bosons ($W^{\pm}/Z$) in a boosted regime. Ideally, the particles originating from the decay of the boosted top lead to the reconstruction of a large-R jet, comprising three-prong substructures, with b- and c-tagged subjets. The SOM algorithm has been demonstrated as a model-agnostic anomaly-finder for probing the rare decay at the LHC, by mapping distinct signal and background regions to separate non-overlapping clusters on the Kohnen map. This helps to identify signal regions with higher signal significances. We also discuss the robustness of this algorithm, especially for other BSM probes with model-agnostic and model-dependent searches.
@@ 1 Introduction
As an illustration, we probed a rare top decay, driven by a flavor changing neutral current (FCNC), to a c-quark and Higgs boson $\left(t\to cH\right)$ with suppressed branching fraction $\lesssim 0.1\%$ [43, 44] in a boosted regime ($p^{top}_{T}>$ 350 GeV), over the background dominated by the most abundant decay mode of the top $\left(t\to bW\right)$, and spurious QCD events $\left(pp\to jj\right)$ mimicking the signal topology. The decay products of the boosted top are highly collimated, enveloped by a large-R jet comprising a three-prong substructure at LHC. Kinematic features specific to the signal $\left(t\to cH\right)$, irreducible background $\left(t\to bW\right)$, and the QCD di-jet background, are mapped on the planar grid which, through iterative adjustment of weights attributed on the nodes, leads to the identification of individual decay modes as clusters appearing on the grids. By clustering the input data, the SOM algorithm effectively guided the determination of optimal signal regions, achieving a signal significance of $\sim 2\sigma$ for $\mathcal{BR}(t\to cH)$ = 0.05% at 3 ab-1 over SM backgrounds, for probing the rare top decay mediated by FCNC. This is a fair leap as compared to our previous endeavor [45], where we presented a simple ML model for the same probe that outperformed existing top-tagging methodologies [40], tailored to probe the said decay. Finally, we provide an interpretation for each cluster boundary formed on the SOM grid, drawing insights into the relevant features underlying the event topologies of the three different processes under consideration.
@@ 2 Self Organizing Maps: The Algorithm
where, $v$ are the coordinates of the said node, $u$ are the coordinates of the BMU, $W_{v}(t)$ is the weight vector of the said node at time step $t$, $X(t)$ is the input event presented at time step $t$, $\eta(t)$ is time-dependent learning rate, and $\phi(u,v,t)$ is the Neighborhood Distance Function for the BMU (at $u$) and the concerned node (at $v$) at the time step $t$. The learning rate $\eta(t)$ follows an exponential damping given as,
@@ 3 Event Generation and Reconstructed Objects
      N-subjettiness ($\tau^{\beta}_{N}$): N-subjettiness characterizes the substructure of boosted jets [61]. Considering N hypothetical subjet axes, N-subjettiness is determined by summing up the angular distances of jet constituents with respect to their respective closest subjet axes. Essentially, it indicates how well a jet aligns with N subjet-structure hypothesis. To establish the prominence of the existence of three subjet axes within the large-R jets of the signal (or the irreducible background) over two, and two subjet axes over one (for QCD di-jet background), we analyze the ratios of N-subjettiness measurements, specifically measuring $\tau_{N}^{\beta}/\tau_{N-1}^{\beta}$, labeled $\tau_{N(N-1)}^{\beta}$, for respectively N=3,2N=3,~2, considering the angular weightage exponent $\beta=1$. These ratios enables the isolation of the QCD multijets from top jet samples.
@@
      Energy- and Mass-Fraction Information: To measure how the energy of a large-R jet is distributed among its subjets, we looked at the fraction of its total energy contained within its constituent subjets. For a given subjet, this is measured as, $\Delta E_{T}^{subjet}=\frac{E_{subjet}}{E_{topjet}}$. We further estimate the composition of large-R jet, with an observable measuring the fractional mass retained after removing the invariant mass of specific subjet combinations (indexed by i,ji,~j): $\Delta X_{ij}=\left(1-\frac{M_{ij}}{M_{topjet}}\right)$. Accurate identification of light jets is crucial for this analysis, and is achieved through a $\chi^{2}$-analysis detailed in [45].
@@ 4.2 Robustness of the cluster boundaries
    where, ‘$a$’ denotes the mean distance between e and all other events within its own cluster, and ‘$b$’ is the mean distance between e and all events in its nearest-neighboring cluster. The average value of $S$ for all events gives the average Silhouette score. This score ranges between [1,1][-1,1], where positive values indicate moderate to good clustering and negative values suggest poor clustering. A score close to zero implies that the events are globally populated in the vicinity of the cluster boundaries.
@@ 4.4 Clustering efficiency with Test dataset
    Test dataset2A/2B/2C: $pp\to tW$ events with top quark decaying through an FCNC decay, $t\to cH$ followed by $H\to b\bar{b}$$H\to WW^{*}$$H\to ZZ^{*}$ decay. The $W$-boson produced in association with the top quark decays leptonically.
@@
    Test dataset3A/3B: The most dominant channel, $pp\to t\bar{t}$, with one of the tops decaying leptonically111 The simulation methodology for the $t\bar{t}$ events, including the software suites, is identical to that described in Section 3. $(t\to bW,W\to\ell\nu_{\ell})$, and the other top undergoes $t\to bW$ decay or $t\to cH$ decay with WW/$H$ decaying hadronically.
@@ 5 BSM Reach: Event Selection to Signal Significances
We now proceed to estimate the reach of the BSM physics process, viz. the FCNC decay of the top quark via $t\to cH$ process, at the high-luminosity run of LHC. To identify the new-physics rich clusters (or, equivalently, the clusters with anomalous events), we first apply a pre-selection cut (trigger requirement): we select events with at least one lepton (electron/muon) with $p_{T}>250$ GeV and missing transverse energy $\cancel{P_{T}}>$ 200 GeV. This choice will ensure that we are indeed focusing on the phase space when the top quarks are produced with sufficient boost.
@@
The first signal region (SR-I) is almost the same as ATLAS and CMS study [65, 44], with the only difference being that the requirements are made at the subjet level. We extend the selection strategy of SR-I to SR-II, accommodating the further requirement of c-tagged jets within the large-R jet. It is needless to say that SR-II is expected to be the most efficient selection strategy in probing the top FCNC decay we are interested in. Finally, we also study the sensitivity of a relatively relaxed criterion to the multiplicity of b-tagged jets within the large R jet by introducing the third signal region, SR-III. Note that, for both SR-I and SR-II, the presence of at least two b-tagged jets enables us to calculate the invariant mass of the two b-tagged jets, and compare the reconstructed mass with the Higgs boson mass. In Figure 7, we show the distribution of the invariant mass mentioned above for the SM Top samples, Test dataset2A, Test dataset3A and Test dataset3B that satisfies the criteria for SR-I and SR-II without the invariant mass cut. For SR-I and SR-II, therefore, we further demand that the invariant mass of the two leading b-tagged jets satisfy the following: $M_{bb}=[110,140]~{\rm GeV}$. The details of the signal regions are listed in Table 4.
@@
Table 4: Signal regions defined to probe the top FCNC decay. Here the observables nSubjetnSubjet, $N_{b-jet}$ and $N_{c-jet}$ denote the total number of subjets, number of b-tagged jets and number of c-tagged jets within the large-R jet, respectively. The invariant mass of the two leading b-tagged jets is denoted by the quantity $M_{bb}$.
As noted previously, the FCNC interactions between the top quark and the Higgs boson via the tcHt-c-H coupling are essential in our quest for BSM physics. To evaluate the feasibility of detecting this exotic mode in the boosted regime using the SOM algorithm, we compute the signal significance, defined as $\mathcal{S}=\sqrt{2\left[(S+B)\ln\left(1+\frac{S}{B}\right)-S\right]}$ [66] where S and B represent the number of signal and background events that meet the jet selection criteria and satisfy the trigger/pre-selection requirement. Before we proceed to calculate the signal significances for the three signal regions, it is important to estimate the cluster efficiencies for each of the SRs. In Table 5, we show the leading order (in $\alpha_{s}$) production cross section for all the processes considered here with $\sqrt{s}=13~{\rm TeV}$. The last column in Table 5 estimates the expected number of events at the 13 TeV run of LHC with integrated luminosity $\mathcal{L}=3000~fb^{-1}$. It is important to note that even though QCD di-jet events seem to overshadow signal events as seen in $3^{rd}$ column of $2^{nd}$ row, the number of events that eventually satisfy the trigger and jet selection cuts, and contribute to the signal regions defined in Table 4 is negligible.
[…]
@@ C-1 C-2 C-3 C-4 C-5 C-6 C-7
where $\sigma$ denotes the cross section of the process, whereas combined branching fraction leading to the final states denoted as $\rm BR$} with $\mathcal{L}$ being the integrated luminosity. The quantity $\epsilon_{cl}$ denotes the number of events passing the pre-selection and trigger requirements for signal region for each cluster over the total number of test samples. The total effective number of events for the signal and backgrounds is labeled as S and B, respectively, with $\rm S=S2A+S2B+S2C+S3B$}, and $\rm B=B1+B2$} where B1 and B2 are the effective number of events corresponding to the two background datasets, i.e., SM top and Test dataset3A.
@@ 6 Summary and Outlook
Probes of processes involving rare decays driven by Flavor-Changing Neutral Current (FCNC) offer a powerful avenue for exploring Physics beyond the Standard Model. The rare decay, $t\to cH$, considered in this article, offers a sensitive probe of potential New Physics interactions, in particular, the tcHt-c-H coupling. In contrast to the supervised learning methodology presented in [45], we present a complementary, model-independent analysis based on the unsupervised learning technique, targeting the boosted regime of this decay. With judicious choices of jet-substructure based observables along with efficient flavor tagging of subjets, similar to the approach followed in [45], we achieve an enhanced signal significance over our previous analysis [45], which outperformed the performances of advanced ML-driven approaches, viz., GNN-based LorentzNet and Transformer-based classifiers.
@@
In [45], we argued that the judicious choice of jet-substructure based observables provided an edge over the GNN-based models which train on the 4-vectors of the large-R jet constituents. With the jet substructures of the signal $\left(t\to cH,~H\to b\bar{b}\right)$, bearing close resemblance to those of the irreducible background $\left(t\to bW,~W\to c\bar{s}\right)$, this schema achieved relatively lower significance. The transformer-based approaches implement an attention mechanism, which determines the relevance of input features for proper identification signals from the background, thereby allowing a dense deep-learning classifier to selectively focus on relevant features suppressing the irrelevant ones, consequently enhancing the accuracy and efficiency of signal recognition. We observed that a simple boosting algorithm with decision tree stumps trained over judiciously chosen jet-substructure-based features, nevertheless, exhibited higher efficiency. This indicates that our choice of input features made the feature space so fragmented that the dominance of a specific class (signal or different SM backgrounds) in a given subspace becomes prominent.
@@
The self-organizing maps, on the contrary, follows a different approach. The nodes, placed at the intersection of its grids, attempt to mimic the feature vector for each input data-point (event topologies) iteratively. The weights are adjusted so that the best matching unit (BMU) and the nodes in its neighborhood have their respective weights aligned with the matched feature vector. With training over each iteration (and updating weights assigned to each node), and subsequently over every batch (and subsequent updation of model parameters), the model captures the latent features specific to signal and different backgrounds. Finally, the grids end up forming clusters, with every node in each cluster possessing a weight vector resembling the underlying features that distinguish different event classes. As discussed in Section 5, each cluster predominantly captures events belonging to a specific Test dataset (see, Tables 6 to 8). The clusters formed on SOM grid are the planar projection of the feature space of the events, in a way that allows events with similar underlying topology captured in a single cluster.
@@
Training SOM with SM top quark and QCD di-jet samples with assessment of topographic error in clustering, we observe that the emerged map fairly preserves the data topography within the grid. Further testing of the trained model with FCNC-driven top decay to a charm quark and a Higgs boson ($t\to cH$) demonstrates remarkable performance in the identification of anomalous regions for the training samples. We define three signal regions (SRs) based on the three-prong substructure of the candidate top jet and their kinematics. For SR-II, our model achieves a significance of $\sim 2\sigma$ for $\mathcal{BR}(t\to cH)$ = 0.05%, establishing the performance of the model for segregating rare FCNC-mediated decays from overwhelming SM backgrounds. We emphasize that a possible explanation behind an enhanced performance of SOM over the existing approaches is embedded in the working principle of the algorithm, which learns to mimic the event features, leading to better capturing of patterns specific to the signal and different backgrounds in distinct clusters. Therefore, SOM has helped to gain insight from the data in an unsupervised manner, specifically by projecting different signal and background regions onto different, non-overlapping clusters within the Kohnen map. This in turn enables the identification of signal regions with high signal significances for probing the FCNC decay of the top quark. As shown in Section 4, SOM has achieved fairly substantial significance despite the vanishingly small $t\to cH$ branching fraction on an overwhelming SM background. It is interesting to note that SOM can also be tailored for supervised learning, upon training the grids with the signal events and the SM backgrounds, thereby estimating the clusters and the event density in each cluster, using SOM as a model-dependent probe of the rare top-decay.
@@
SOMs, since its inception, has outgrown as tool for clustering and dimensionality reduction. As mentioned above, it has been adapted for supervised learning, enabling enhanced model performance while simultaneously preserving the topography of the training dataset. This versatility makes SOM a potential tool for BSM probes. Further improvements can be explored with alternative distance metrics, viz., the Mahalanobis distance, or the cosine similarity, in order to obtain more distinct clusters of signals and backgrounds in the Kohonen plane. Adaptive frameworks, such as Growing Self-Organizing Map (GSOM), which dynamically adapts its structure during training, by adding nodes as required based on data complexity, could further optimize the signal-to-background ratio. SOMs and its extensions are therefore, potential tools for future phenomenological explorations.